a researcher reports that the average salary of a teacher php 27000 a sample of 32 teachers has a mean of php 29500 at a=0.01 test the claim that the co teachers earn more than php 27000 a months the standard deviation of the population is php 2 250

1
Expert's answer
2022-06-08T13:18:58-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le27000"

"H_1:\mu>27000"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z:z>2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{29500-27000}{2250/\sqrt{32}}=6.2854"

Since it is observed that "z=6.2854>2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z>6.2854)= 0," and since "p= 0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 27000, at the "\alpha = 0.01" significance level.